OpenStudy (anonymous):

s=2(lw+hw+hl) solve for W

OpenStudy (anonymous):

1. distribute the 2 2. subtract 2hl from both sides 3. factor a w to get w(2l +2h) you should have the following \[s - 2hl= w(2l+2h) \] divided both sides 2l + 2h to get w.

OpenStudy (anonymous):

Yeah I have gotten there and I type as my answer W = s-2hl/2l+2h and am getting the wrong answer

OpenStudy (anonymous):

try adding parentheses around the top term and the bottom term if you're using an online homework application and it's not entering in correctly

OpenStudy (anonymous):

I have, and even simplified the bottom as 2(l+h) maybe I am missing another simplification?

OpenStudy (anonymous):

hmm there isn't any other simplification that I can see... it could just be a glitch in the program or maybe you capitalized something and it didn't recognize it because of that? these are really touchy sometimes.

OpenStudy (anonymous):

make sure you use parenthesis properly. otherwise, your answer could be read as wrong. use \[(s-2hl)/(2l + 2h)\]